OpenStudy (anonymous):

Conside the curve with parametric equations: x=4t^2-5, y=2t+3 dy/dx, the equation of the tangent line when t=2, and the point(s) where there is a vertical tangent line

OpenStudy (anonymous):

dy/dx=2/8t = 1/4t

OpenStudy (anonymous):

do we need coordinates to find the equation of the tangent line or do we plug the value of t into the original x and y equations to find the x and y coordinates

OpenStudy (anonymous):

You can plug in t. Thats the point of generalizing a derivative to handle parametrically defined curves.

OpenStudy (anonymous):

yes i understand that part now but how do i find the points where there is a vertical line when y=0?

OpenStudy (anonymous):

You mean, where is there a vertical line? Wherever the denominator is undefined. Namely, t=0. So the point would be: (-5,3)

OpenStudy (anonymous):

i dont see how you got that...

OpenStudy (anonymous):

oh so we plug 0 into the original equations.....okay sorry slow moment THANKS MALEVOLENCE!!

OpenStudy (anonymous):

Haha, its okay :P No problem xD